OpenStudy (anonymous):

the difference between two numbers is 4. three times the smaller number minus four times the larger is -28. what are the numbers

OpenStudy (anonymous):

let two numbers be x & y x - y = 4 and 3y - 4x = -28 multiplying first equation by 3 and adding it to second eqn we get 3y - 4x + ( 3x - 3y ) = -28 + 12 - x = -16 x = 16 y = 12

OpenStudy (anonymous):

sorry, does not line up

jhonyy9 (jhonyy9):

smaller s and larger l so than l-s=4 3s-4l= -28 - so multiply first equation by 3 and assume than will get 3l-3s=12 3s-4l= -28 ------------ 3l+(-4l)+(-3s)+3s= -16 -l+0=-16 l=16 so than we know that l-s=4 so than 16-s=4 so than s=12 so l=16 s=12